Hawk Host Status

Washington, DC

Washington DC Datacenter

Time Status Message
8 months, 3 weeks ago The wdc004 system is now back online
8 months, 3 weeks ago The wdc004 system has gone unresponsive again suggesting an underlying hardware issue. We're currently investigating
8 months, 3 weeks ago The wdc004 system is now back online
8 months, 3 weeks ago The wdc004 system has had a kernel panic. We're currently investigating
9 months, 3 weeks ago We swapped out all of the RAM for WDC004.HAWKHOST.COM and the machine is online at this time.
9 months, 3 weeks ago The shared hosting server WDC004.HAWKHOST.COM is experiencing hardware issues. We're having technicians in the data center investigate.
9 months, 4 weeks ago The system is now back online
9 months, 4 weeks ago The wdc-sd010 system requires a file system check. We are performing one now.
11 months ago The wdc-vz004 system is now back online
11 months ago The wdc-vz004 system is being taken down for emergency maintenance. We will update this when the system is back online
11 months ago The wdc-vz003 system is now back online
11 months ago The wdc-vz003 system is being taken down for emergency maintenance. We will update this when the system is back online
1 year, 1 month ago The attack against wdc011 has now been filtered.
1 year, 1 month ago There is a large scale distributed denial of service attack targeting the wdc011 system. It is causing a large amount of packet loss as a result. We're currently working on filtering the attack.
1 year, 2 months ago The wdc012 system is now back online.
1 year, 2 months ago The wdc012 system is currently down due to bad memory. Our team are currently working on replacing the memory on the system.
1 year, 9 months ago At this time full network redundancy has been restored and all servers are online.
1 year, 9 months ago Network engineers have corrected the boot issues on the front-end public network router fcr04a.wdc01 and it is online at this time, public network should again be accessible for customers. The redundant router is currently booting, and we expect full network capacity and redundancy to be restored within 5 minutes.
1 year, 9 months ago The front-end public router experienced a trouble during its boot up, and network engineers are working on correcting that. At this time the front-end public network is still experiencing intermittent connectivity issues while engineers work on the issue.
1 year, 9 months ago The majority of systems are now back online. Our team will update individual system notices for machines that did not come back up properly after power was restored.
1 year, 9 months ago At 20-Jul-2015 07:48 UTC, the facility provider has been successful in restoring power to Pod04 of WDC01. At this time, the routers are coming online and datacenter site staff are bringing all servers back online. This process may take up to 15 minutes.
1 year, 9 months ago Our facility provider has informed us that the vendor of the power system is on-site, and that they have attempted twice thus far to transfer power over to generator. Unfortunately, a continued issue with the power switching system has not allowed the transfer to remain stable. As a result a portion of our Washington DC location continue to be without power.
1 year, 9 months ago Our facilities provider has confirmed that there was a utility power disruption earlier in the morning, and that the automatic transfer switch (ATS) failed to transfer power to generator backup. At this time facilities engineers are working on troubleshooting the ATS trouble so that they can affect the switchover. At this time servers affected in Washington DC continue to be without power.
1 year, 9 months ago At 20-Jul-2015 11:04 UTC Server Room 02 of the Washington DC datacenter experienced power disruption. Datacenter engineers are currently working on restoring power.
2 years, 8 months ago At approximately 4:04am local time the fiber splicing was completed and complete network connectivity was restored.
2 years, 8 months ago We have received an update from the provider that Fiber crews will need to replace 2500 - 3000 ft of cable. The new fiber reel is en route and the crews are expecting to have the new cable spliced by 2014-08-27 2:00 UTC.
2 years, 8 months ago Multiple redundant links between the Datacenter Aggregation Routers (DAR) in WDC01 and the Backbone routers in WDC02 (BBR) are offline. The Networking team is working with the provider to restore service to the affected links as quickly as possible. However, the provider has identified that a fiber cut is responsible and that technicians have been dispatched to correct the issue. Due to the reduction in overall network capacity between WDC01 and backbone network, while these links are out of service customers may experience higher than normal network latency and some packet loss.